OpenStudy (narissa):

I need help showing my work---- Suppose you paid 88.00 for a pair of shoes tht has been discounted 20%. What was the original price of the shoes before the discount? a)identify your unknown variable b) set up a porportion that represents the problem c) solve the porportion,showing all work! d) clearly state your answer in complete sentence

OpenStudy (narissa):

@HossamHoussien

OpenStudy (narissa):

do u know how to do this @HossamHoussien

OpenStudy (narissa):

would a) be something like 88-20%=x

OpenStudy (hossamhoussien):

assume that price before discount was (x) $ now calculate the price after discount with next equation: \[x-\frac{ x \times 20 }{ 100}=88\] \[100x-20x=8800\] \[80x=8800\] \[x=110\]

OpenStudy (narissa):

so thats how i do it or thats an example?

OpenStudy (narissa):

the second part shouldn't it say 80?

OpenStudy (narissa):

im confused @HossamHoussien

OpenStudy (narissa):

how does x=110

OpenStudy (narissa):

so u did it as a b c d

OpenStudy (hossamhoussien):

this is the answer to check if it true or not : price before discount = 110 discount = 20% discount amount = (110*20)/100=22 price after discount = 110-22 = 88

OpenStudy (narissa):

ok so identify the unknown variable i write what u posted first

OpenStudy (hossamhoussien):

a,b,c and d are steps to solve it not options, still confused?

OpenStudy (narissa):

shouldnt the second one be 80 not 8800?

OpenStudy (narissa):

ok so u did it step by step right

OpenStudy (hossamhoussien):

what i posted first was a,b,c and d . bcz I've done all the steps

OpenStudy (narissa):

100x-20x=80 right not 8800?

OpenStudy (hossamhoussien):

yea, and bcz u did multiply all equation terms by 100 so 88*100=8800

OpenStudy (narissa):

ok thanks man i was confused

OpenStudy (hossamhoussien):

i hope you got it ;)

OpenStudy (narissa):

how do i write the last one in a complete sentence the answer

OpenStudy (narissa):

nvm i know the original price of the shoes beefore the discount was 110.00

OpenStudy (narissa):

thanks alot